Explain how you would use monetary policy to reduce unemployment.

What is the investment multiplier and how does it work on the spending in an economy?
Answer #1

Monetary policy lowers rate of interest which reduces the cost of borrowing and gives encouragement to individuals to invest.As a result aggregate demand increases and GDP growth goes up.So demand deficient unemployment gets reduced.Lower interest rates brings down exchange rates and as a result exports increase.Thus monetary policy reduces interest rate to give an impetus to aggregate demand.Lowering of interest rate may not always increase aggregate demand . So central government may increase the money supply known as quantitative easing , to increase demand and boost the economy.

Investment multiplier(k) is the ratio of increase in national income (change in income )as a result of increase in investment(change in investment).

When there is injection of extra income in the economy , there is extra spending which creates more income.The multiplier effect refers to the final increase in income from spending.The size of the multiplier is related to the households marginal propensity to consume or marginal propensity to save.It is stated that one person's spending becomes another person's income.AS for eg if 80%of new income is spent , MPC will be 80/100 ie 0.8.So the formula is 1/1-MPC. So if consumers consume 0.8 and save 0.2 of every dollar ,the multiplier will be 1/1-0.8=1/0.2=5